When a firm uses different factors of production or at least cost combination the optimum combination of factors is achieved when MPL / PL = MPL / PL. Intuitively explain this statement


1
Expert's answer
2022-03-08T08:50:48-0500

If the marginal benefit of additional labor, "\frac{MP_{L} }{P_{L}}", is greater than marginal cost, "\frac{MP_{K} }{P_{K}}", then the organization will be better off by spending more on labor and less on capital. Also, if "\frac{MP_{K} }{P_{K}}" is exceeds "\frac{MP_{L} }{P_{L}}" , the firm should be spending more on capital and less on labor. When a business generates the maximum quantity of output for a given cost, it is said to be in equilibrium given by: "\frac{MP_{L} }{P_{L}} = \frac{MP_{K} }{P_{K}}".